OpenStudy (anonymous):

x^2-7x+11 find x and y intercepts

OpenStudy (anonymous):

i got (3.5,0) and (0,11) is this right?

OpenStudy (anonymous):

x intercepts : x1,2 = (7 +/- (49 - 44) ^0.5) / 2 x1 = (7 + 5^0.5) / 2 x2 = (7 -5^0.5) / 2 y intercepts : plug 0 in x then get y =11

hero (hero):

I was imagining something else.